OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (aum):

\(\le\) is called "less than or equal to". \(\ge\) is called "greater than or equal to". What is " less than or greater to" ?

OpenStudy (anonymous):

idk thats how it came out when i copied and pasted should be \[y \le x-5\] \[y \le -x-4\]

OpenStudy (aum):

First graph the two equations: y = x - 5 and y = -x - 4. Then shade the appropriate region that satisfies each inequality. The solution set is the common shaded area. Test each answer to see which point falls in the common shaded area.

OpenStudy (anonymous):



OpenStudy (aum):

Good. The common area shared by both inequalities is: